# JWST PRIMER: A lack of outshining in four normal galaxies from the ALMA-CRISTAL Survey

### v1

@@ 6 A comparison to the FirstLight simulation
In this section we compare our results to the zoom-in simulation, FirstLightĀ (Ceverino etĀ al., 2017). Nakazato etĀ al. (2024) use the FirstLight simulations to provide a spatially resolved analysis of high-redshift galaxies. In particular, they simulated galaxies of comparable stellar masses to the objects we study in this work allowing for a close comparison with ourĀ JWST results. Nakazato etĀ al. (2024) identify 100 parsec scale clumps within their simulated galaxies, similar to the morphology and sizes of clumps that we identify in the CRISTAL galaxies. They found that between 10ā20 percent of galaxies with Ā were observed in a phase dominated by major-merger activity, leading to the observed clump formation, in the redshift range of their work ( to ). While this fraction is lower than we recover, the difference could be due to the definition of clumps inĀ Nakazato etĀ al. (2024) who used a threshold in star-formation surface density on the native resolution simulations. This merger driven clump formation within high-redshift galaxies has also been found in other simulations, such asĀ Ocvirk etĀ al. (2024),Ā Ma etĀ al. (2019) andĀ Pallottini etĀ al. (2022).
@@
Interestingly, the zoom in cosmological simulations are able to recover galaxies with similar extent and clumpy morphology to the CRISTAL galaxies. FigureĀ 8 shows three galaxies from the FirstLight simulation at , with stellar masses and SFRs that closely match those in our sample (ā; SFR ā/yr). The SFRs shown are calculated with stars younger than 10 Myrs. The broad morphology of old (> 100 Myr) and young (< 10 Myr) stars are consistent in the simulated sources, with both the old and young stars showing clumps separated by - kpc within an extended structure. This is very similar to what we observe for the CRISTAL galaxies and hence we can draw parallels with the formation scenario of the simulated galaxies and our observed galaxies. Ā Nakazato etĀ al. (2024) identify two broad populations of clumps, with the majority being formed as a result of major mergers in the simulation. The first class of clumps were found to be older (ages ) and represent proto-bulges, with the second class being younger clumps that have high sSFRs of and are formed within the tidal tails of the interacting systems. We recover both older regions of several tens of Myr, in addition to younger regions that show comparable sSFRs to the simulations (e.g. Fig.Ā 2). In addition, the simulated images show the importance of dust emission on the observed morphology, with the compact cores in the simulated galaxies becoming more diffuse when dust is applied. Our analysis of the ALMA-CRISTAL dust continuum data in SectionĀ 5.4 shows how dust is likely influencing the observed morphology of the most massive galaxies we study here, by obscuring certain regions that appear fainter and redder in the rest-frame UV than other parts of the galaxy.
@@
Overall, the similarity of the CRISTAL galaxies we study to the sources found in the cosmological simulation presented inĀ Nakazato etĀ al. (2024) allows us to tentatively conclude that the multi-wavelength morphology of our sample could have been impacted by major mergers. This conclusion is supported by the initial morpho-kinematic classes of these sources defined inĀ Le FĆØvre etĀ al. (2020) from the ALPINE survey, which indicated dispersion dominated sources. It is now supported by the first kinematic results from the CRISTAL survey, which also find evidence for major mergers in the [CII]Ā distributionĀ (Posses etĀ al., 2024; Solimano etĀ al., 2024). A detailed analysis of the full [CII]Ā kinematics from the CRISTAL program (Lee et al. in prep) will shed more light onto the formation mechanism of galaxies in the early Universe. A caveat of this conclusion is that in the FirstLight simulation only 10ā20 percent of sources show the clumpy morphology that we see in all four of the CRISTAL galaxies in this study. One reason for this could be the different clump selection procedures between the simulation and observational analyses, and further investigation is required to match the population statistics and small scale morphological measurements in models and data (e.g. see the detailed size-luminosity comparison inĀ Varadaraj etĀ al., 2024).
@@ 7 Conclusions
In this work we perform a pixel-by-pixel analysis of four normal galaxies (with ) at ā using the multi-band NIRCam imaging from theĀ JWST PRIMER program. The four sources were previously spectroscopically confirmed as part of the ALPINE program, and have deep, high-spatial resolution, dust continuum observations from the ALMA-CRISTAL large program. Using an SED fitting analysis with a simple constant SFH model we study the resolved and integrated properties and investigate the potential biases present in fitting to unresolved photometry of galaxies of this mass. Our main conclusions are as follows:
ā¢
@@
Compared to the previously availableĀ HST optical and NIR data, we find that all of the sources break into further kpc-scale clumps and show at least 2 (and up to 8) separate components in the high-resolutionĀ JWST data. We find comparable morphologies in the rest-frame UV and optical, with no clear evidence for underlying older disk structures. The non-parametric sizes of the sources are as expected from the size-mass relation, however the full extent of the galaxies is ā taking into account the widely separated clumps.
ā¢
@@
In contrast to some studies of lower mass galaxies, including the strongly lensed Cosmic Grapes source, we do not find extremely young ages (or high rest-frame optical emission lines) or other evidence for strong outshining of the observed global emission by the youngest regions. We find that the stellar masses derived from the resolved pixel-by-pixel analysis only slightly higher (ā) than those determined from fitting to the global integrated photometry. Our results agree with those found in star-forming galaxies of similar mass at Ā (Shen etĀ al., 2024) and Ā (PĆ©rez-GonzĆ”lez etĀ al., 2023). The result holds for both an assumed constant and delayed star-formation history.
ā¢
@@
In comparison to the previously published best-fitting properties of these galaxies, ourĀ JWST analysis recovers younger ages () and bluer rest-frame UV slopes (). TheĀ Ā photometry from the COSMOS2020 catalogue was found to be overestimated by a factor of for CRISTAL-17, likely due to confusion in the low resolutionĀ Spitzer/IRAC data. We further recover slightly weaker HĀ emission than the results obtained in the ALPINE catalogues (Ć ).
ā¢
@@
For two of the sources we compare the physical parameter maps to the resolved dust continuum measurements from the ALMA-CRISTAL program. In both cases we see a strong correlation between the direct detection of dust and the dust attenuation derived from theĀ JWST photometry in the rest-frame UV and optical, as parametrised by the rest-frame UV slope and . The peak of the dust continuum also correlates with the centre of mass of the galaxies, as predicted by the recent simulation work ofĀ Ocvirk etĀ al. (2024).
ā¢
@@
We compute the resolved āĀ relation for CRISTAL-11 and CRISTAL-13 to decouple the effect of geometrically separate regions on the correlation. We find bluer rest-frame UV slopes than that determined in the ALPINE program (from COSMOS2015 photometry). With the assumed dust temperature and parameters fromĀ Mitsuhashi etĀ al. (2023), we recover a correlation between Ā and colour that is in good agreement with the local starburstĀ Calzetti etĀ al. (2000) relation. As expected, the position of the ALMA-CRISTAL dust detection is co-spatially with the redder regions of the galaxies, suggesting that dust attenuation is shaping the observed rest-frame UV morphology.
ā¢
@@
These results demonstrate the power of combiningĀ JWST resolved photometry with high-spatial resolution ALMA imaging to understand galaxies at high redshift. Additional insight into the CRISTAL galaxies will be obtained with upcoming NIRSpec IFU spectra (PI: Faisst, PID: 3045 and PI: Aravena, PID: 5974), which will allow detailed measurements of the stellar ages and ioniziation conditions from the rest-frame optical emission line ratios.
@@ Appendix A Comparison to Li+2024
In TableĀ 4 we present the resolved and integrated Ā and SFRs from the fitting analysis of Li et al. (2024). In Li et al. (2024) they performed SED fitting using MAGPHYS, of a larger sample of CRISTAL galaxies using the full rest-frame UV to FIR data including that fromĀ JWST and the ALMA-CRISTAL program. They matched the spatial resolution of theĀ JWST imaging to that of the available ALMA data, and hence this provides a complementary fitting procedure (at slightly lower spatial resolution) than ourĀ JWST-only analysis. Reassuringly we find excellent agreement between the stellar masses, with the resolved values being within . While the integrated value agrees well for CRISTAL-11, we find a slightly lower (by 0.25ādex) value for CRISTAL-13 although this is only discrepant. The SFRs are all consistent within the errors. The resolved SFRs are slightly lower in our analysis, which could be due to not including the ALMA FIR data within our SED fitting. The difference is only however, demonstrating that in these sources the fitting ofĀ JWST only photometry is able to recover over 80 percent of the total SFR (as defined as that from the full rest-frame UV to FIR fitting). This result is consistent with our āĀ analysis, which shows that first, the galaxies are relatively blue, suggestive of a moderate impact of dust obscuration on the observed rest-frame UV light. And second, that theĀ Calzetti etĀ al. (2000) law provides a good description of theĀ āĀ relationship. We fit assuming this dust law in our BAGPIPES fitting, and this appears to account for the majority of obscured SFR when fitting to the rest-frame UV and optical light alone.
@@ 4 Department of Physics, The University of Tokyo, 7-3-1 Hongo, Bunkyo, Tokyo 113-0033, Japan
5Instituto de Estudios AstrofĆsicos, Facultad de IngenierĆa y Ciencias, Universidad Diego Portales, Av. EjĆ©rcito Libertador 441, Santiago, Chile
@@ 8 CIAFF, Facultad de Ciencias, Universidad Autonoma de Madrid, E-28049 Madrid, Spain
15Departamento de AstronomĆa, Universidad de Concepción, Barrio Universitario, Concepción, Chile
@@
20Centro de AstrobiologĆa (CAB), CSIC-INTA, Ctra. de Ajalvir km 4, Torrejón de Ardoz, E-28850, Madrid, Spain

### v2

@@ 6 A comparison to the FirstLight simulation
In this section we compare our results to the zoom-in simulation, FirstLight (Ceverino et al., 2017). Nakazato et al. (2024) use the FirstLight simulations to provide a spatially resolved analysis of high-redshift galaxies. In particular, they simulated galaxies of comparable stellar masses to the objects we study in this work allowing for a close comparison with our JWST results. Nakazato et al. (2024) identify 100 parsec scale clumps within their simulated galaxies, similar to the morphology and sizes of clumps that we identify in the CRISTAL galaxies. They found that between 10–20 percent of galaxies with were observed in a phase dominated by major-merger activity, leading to the observed clump formation, in the redshift range of their work ( to ). While this fraction is lower than we recover, the difference could be due to the definition of clumps in Nakazato et al. (2024) who used a threshold in star-formation surface density on the native resolution simulations. This merger driven clump formation within high-redshift galaxies has also been found in other simulations, such as Ocvirk et al. (2024), Ma et al. (2019) and Pallottini et al. (2022).
@@
Interestingly, the zoom in cosmological simulations are able to recover galaxies with similar extent and clumpy morphology to the CRISTAL galaxies. Figure 8 shows three galaxies from the FirstLight simulation at , with stellar masses and SFRs that closely match those in our sample (–; SFR –/yr). The SFRs shown are calculated with stars younger than 10 Myrs. The broad morphology of old (> 100 Myr) and young (< 10 Myr) stars are consistent in the simulated sources, with both the old and young stars showing clumps separated by - kpc within an extended structure. This is very similar to what we observe for the CRISTAL galaxies and hence we can draw parallels with the formation scenario of the simulated galaxies and our observed galaxies. Nakazato et al. (2024) identify two broad populations of clumps, with the majority being formed as a result of major mergers in the simulation. The first class of clumps were found to be older (ages ) and represent proto-bulges, with the second class being younger clumps that have high sSFRs of and are formed within the tidal tails of the interacting systems. We recover both older regions of several tens of Myr, in addition to younger regions that show comparable sSFRs to the simulations (e.g. Fig. 2). In addition, the simulated images show the importance of dust emission on the observed morphology, with the compact cores in the simulated galaxies becoming more diffuse when dust is applied. Our analysis of the ALMA-CRISTAL dust continuum data in Section 5.4 shows how dust is likely influencing the observed morphology of the most massive galaxies we study here, by obscuring certain regions that appear fainter and redder in the rest-frame UV than other parts of the galaxy.
@@
Overall, the similarity of the CRISTAL galaxies we study to the sources found in the cosmological simulation presented in Nakazato et al. (2024) allows us to tentatively conclude that the multi-wavelength morphology of our sample could have been impacted by major mergers. This conclusion is supported by the initial morpho-kinematic classes of these sources defined in Le Fèvre et al. (2020) from the ALPINE survey, which indicated dispersion dominated sources. It is now supported by the first kinematic results from the CRISTAL survey, which also find evidence for major mergers in the [CII] distribution (Posses et al., 2024; Solimano et al., 2024). A detailed analysis of the full [CII] kinematics from the CRISTAL program (Lee et al. in prep) will shed more light onto the formation mechanism of galaxies in the early Universe. A caveat of this conclusion is that in the FirstLight simulation only 10–20 percent of sources show the clumpy morphology that we see in all four of the CRISTAL galaxies in this study. One reason for this could be the different clump selection procedures between the simulation and observational analyses, and further investigation is required to match the population statistics and small scale morphological measurements in models and data (e.g. see the detailed size-luminosity comparison in Varadaraj et al., 2024).
@@ 7 Conclusions
In this work we perform a pixel-by-pixel analysis of four normal galaxies (with ) at – using the multi-band NIRCam imaging from the JWST PRIMER program. The four sources were previously spectroscopically confirmed as part of the ALPINE program, and have deep, high-spatial resolution, dust continuum observations from the ALMA-CRISTAL large program. Using an SED fitting analysis with a simple constant SFH model we study the resolved and integrated properties and investigate the potential biases present in fitting to unresolved photometry of galaxies of this mass. Our main conclusions are as follows:
@@
Compared to the previously available HST optical and NIR data, we find that all of the sources break into further kpc-scale clumps and show at least 2 (and up to 8) separate components in the high-resolution JWST data. We find comparable morphologies in the rest-frame UV and optical, with no clear evidence for underlying older disk structures. The non-parametric sizes of the sources are as expected from the size-mass relation, however the full extent of the galaxies is – taking into account the widely separated clumps.
@@
In contrast to some studies of lower mass galaxies, including the strongly lensed Cosmic Grapes source, we do not find extremely young ages (or high rest-frame optical emission lines) or other evidence for strong outshining of the observed global emission by the youngest regions. We find that the stellar masses derived from the resolved pixel-by-pixel analysis only slightly higher (–) than those determined from fitting to the global integrated photometry. Our results agree with those found in star-forming galaxies of similar mass at (Shen et al., 2024) and (Pérez-González et al., 2023). The result holds for both an assumed constant and delayed star-formation history.
@@
In comparison to the previously published best-fitting properties of these galaxies, our JWST analysis recovers younger ages () and bluer rest-frame UV slopes (). The photometry from the COSMOS2020 catalogue was found to be overestimated by a factor of for CRISTAL-17, likely due to confusion in the low resolution Spitzer/IRAC data. We further recover slightly weaker H emission than the results obtained in the ALPINE catalogues (Å).
@@
For two of the sources we compare the physical parameter maps to the resolved dust continuum measurements from the ALMA-CRISTAL program. In both cases we see a strong correlation between the direct detection of dust and the dust attenuation derived from the JWST photometry in the rest-frame UV and optical, as parametrised by the rest-frame UV slope and . The peak of the dust continuum also correlates with the centre of mass of the galaxies, as predicted by the recent simulation work of Ocvirk et al. (2024).
@@
We compute the resolved – relation for CRISTAL-11 and CRISTAL-13 to decouple the effect of geometrically separate regions on the correlation. We find bluer rest-frame UV slopes than that determined in the ALPINE program (from COSMOS2015 photometry). With the assumed dust temperature and parameters from Mitsuhashi et al. (2023), we recover a correlation between and colour that is in good agreement with the local starburst Calzetti et al. (2000) relation. As expected, the position of the ALMA-CRISTAL dust detection is co-spatially with the redder regions of the galaxies, suggesting that dust attenuation is shaping the observed rest-frame UV morphology.
@@
These results demonstrate the power of combining JWST resolved photometry with high-spatial resolution ALMA imaging to understand galaxies at high redshift. Additional insight into the CRISTAL galaxies will be obtained with upcoming NIRSpec IFU spectra (PI: Faisst, PID: 3045 and PI: Aravena, PID: 5974), which will allow detailed measurements of the stellar ages and ioniziation conditions from the rest-frame optical emission line ratios.
@@ Appendix A Comparison to Li+2024
In Table 4 we present the resolved and integrated and SFRs from the fitting analysis of Li et al. (2024). In Li et al. (2024) they performed SED fitting using MAGPHYS, of a larger sample of CRISTAL galaxies using the full rest-frame UV to FIR data including that from JWST and the ALMA-CRISTAL program. They matched the spatial resolution of the JWST imaging to that of the available ALMA data, and hence this provides a complementary fitting procedure (at slightly lower spatial resolution) than our JWST-only analysis. Reassuringly we find excellent agreement between the stellar masses, with the resolved values being within . While the integrated value agrees well for CRISTAL-11, we find a slightly lower (by 0.25 dex) value for CRISTAL-13 although this is only discrepant. The SFRs are all consistent within the errors. The resolved SFRs are slightly lower in our analysis, which could be due to not including the ALMA FIR data within our SED fitting. The difference is only however, demonstrating that in these sources the fitting of JWST only photometry is able to recover over 80 percent of the total SFR (as defined as that from the full rest-frame UV to FIR fitting). This result is consistent with our – analysis, which shows that first, the galaxies are relatively blue, suggestive of a moderate impact of dust obscuration on the observed rest-frame UV light. And second, that the Calzetti et al. (2000) law provides a good description of the – relationship. We fit assuming this dust law in our BAGPIPES fitting, and this appears to account for the majority of obscured SFR when fitting to the rest-frame UV and optical light alone.
@@ Appendix B Delayed exponential SFH
Appendix C Corner plots for pixel fits
Here we present corner plots as derived from BAGPIPES for individual pixels in our resolved fitting. The relevant pixels, chosen to represent regions of the galaxy with different derived properties, are highlighted in Fig. 2 and Fig. 3. We recover well known degeneracies in our fitting, in particular the degeneracy between age and stellar mass, and to a lesser degree dust and stellar mass. The coverage of strong rest-frame optical emission lines reduces the effect of these degeneracies, as the strength of the H and H + [OIII] lines we infer from our photometry constrains the age in the individual pixel (and integrated) fits. In some cases we hit the edge of parameter space, in particular for stellar age. Demonstrating the likely presence of very young stellar populations within the sources, as expected for star-forming regions in UV-bright LBGs.
@@ 4 Department of Physics, The University of Tokyo, 7-3-1 Hongo, Bunkyo, Tokyo 113-0033, Japan
5Instituto de Estudios Astrofísicos, Facultad de Ingeniería y Ciencias, Universidad Diego Portales, Av. Ejército Libertador 441, Santiago, Chile
@@ 8 CIAFF, Facultad de Ciencias, Universidad Autonoma de Madrid, E-28049 Madrid, Spain
15Departamento de Astronomía, Universidad de Concepción, Barrio Universitario, Concepción, Chile
@@
20Centro de Astrobiología (CAB), CSIC-INTA, Ctra. de Ajalvir km 4, Torrejón de Ardoz, E-28850, Madrid, Spain